Good Eats, Inc., a national food company, is evaluating whether to extend its product line into plant-based foods. Given the recent success of Beyond Meat’s IPO and Impossible Foods’ anticipated IPO, Good Eats believes it can acquire the R&D talent to quickly develop similar plant-based products. The new product line would be marketed through its current distribution channels.

Good Eats’ management will introduce the new product line early next year. The CFO believes the plant-based product line will require $40 million in additional net operating assets for the R&D and related production facilities as well as supporting the sales initiative. The CFO estimates that the new plant-based product line sales will be $20 million.

For the current year, Good Eats anticipates its existing business will earn $480 million in sales generating a profit margin of 3.5 percent with an asset turnover of 3.0. When the new product line comes on board next year, the existing business line sales will be $480 million as the new product line will cannibalize some sales initially. However, Good Eats’ expected profit margin will increase to 4.3 percent on all sales.

Future growth for the plant-based food line is expected to be greater than the traditional food line, although management is uncertain what the difference will be as other new entrants will certainly enter into the plant-based food market as well.

Based solely on the above financial consideration, explain whether you would recommend management consider extending its product line into plant-based foods. [Hint: Conduct a RNOA Analysis]
Answer #1

The credibility of launching the plant-based foods can be found out by computing the Return on Net Operating Assets (RNOA).

RNOA = net income / assets(fixed assets and working capital)

Without the proposed plan:

Asset turnover = 3
net sales/ average total assets = 3
480mn/assets = 3
Thus, assets = 480/3 = 160mn

Profit margin = 3.5% of sales
profit = 3.5% * 480 = 16.8mn

Thus, RNOA = 16.8/160 = 0.105 of 10.5%

With the proposed plan:

Total Assets required = 160+40 = 200mn

Total sales estimated = 480+20 = 500mn

Profit margin = 4.3%
= 4.3% * 500
= 21.5mn

Thus RNOA = 21.5 / 200 = 0.1075 or 10.75%

Since the RNOA for the proposed plan is better, it is advisable to expand the business with the plant based foods.